Please show any formulas and steps

1) Some laboratory equipment sells for $85,000. The manufacturer offers financing at 6% with annual payments for 5 years for up to $65,000 of the cost. The salesman is willing to cut the price by 10% if you pay cash. What is the interest rate you would pay by financing?
Answer #1

The interest rate is computed as shown below:

$ 65,000 x (1 + r / 12) 5 x 12 = ($ 85,000 - 10% x $ 85,000) x (1 + 0.06)5

$ 65,000 x (1 + r / 12)60 = $ 76,500 x 1.338225578

$ 65,000 x (1 + r / 12)60 = $ 102,374.2567

(1 + r / 12)60 = $ 102,374.2567 / $ 65,000

(1 + r / 12)60 = $ 1.574988564

r / 12 = ($ 1.574988564) 1 / 60 - 1

r = 9.12% Approximately

Feel free to ask in case of any query relating to this question